# Conditional regression for the Nonlinear Single-Variable Model

### v1

@@ C.0.1 Background: standard & modified Meyer’s staircase
    $\begin{split}I\to L^{2}(I)\quad:\quad\mu\mapsto g_{\mu}(t):=\frac{1}{\sqrt[4]{2\pi}\sqrt{\delta}}\exp\left(-\frac{|t-\mu|^{2}}{4\delta^{2}}\right)\end{split}$ (10)
@@
    $\begin{split}L^{2}(I)\to\mathbb{R}^{d}\quad:\quad g_{\mu}(t)\mapsto(g_{\mu}(1/d),\dots,g_{\mu}(1))^{\intercal}\end{split}$ (11)
@@
    $x_{k}(t)=\frac{1}{\sqrt[4]{2\pi}\sqrt{\delta}}\exp\left(-\frac{|k/d-t|^{2}}{4\delta^{2}}\right)$ (12)
@@
    (ii) the local reach of modified Meyer’s staircase is almost the reciprocal of the magnitude of local curvature. The above two aspects indicate that the curve traverses the space with weak self-entanglement. Hence, it suggests the possibility of finding a linear projection $P:\mathbb{R}^{d}\to\mathbb{R}^{d^{\prime}}$ with $d^{\prime}$ much smaller than $d$ such that the projected image $P\gamma$ is a much simpler curve. For example, suppose we perform the linear projection of modified Meyer’s staircase onto its first few principal components: the projected image is a simple curve, and the learning problem can be significantly simplified if we study the regression problem on the projected curve. Hence, to test the performance algorithm for the regression problem, we aim to test curves that are so complex that there is no trivial dimension reduction, e.g. via standard techniques such as Principal Component Analysis.
@@ C.0.2 Meyer helix
    $x_{k}(t)=\frac{1}{\sqrt[4]{2\pi}\sqrt{\delta_{d}}}\cos\left(ak+\frac{t-k/d}{\delta_{d}^{\prime}}\right)G\left(\frac{\left|k/d-t\right|}{\delta_{d}}\right)$ (14)
@@
    Moreover, the above properties of complexity are pretty robust: we can also choose Gaussian-type to decay $G(z)=\exp(-z^{2})$ or choose another constant in the variants (e.g.,, $a,\delta_{d}$, and $\delta_{d}^{\prime}$) and the above properties persist. This indicates that this collection of Meyer helix curves indeed has its complexity growing with ambient dimension $d$. Numerical tests suggest that this collection of curves does not have simple dimension reduction via random projections that preserves geometric properties such that length and reach. Recall the following form of Johnson-Lindenstrauss random projection lemma for manifolds:
@@
    $(1-\epsilon)\sqrt{\frac{M}{N}}\leqslant\frac{\left\|\Phi x-\Phi y\right\|_{2}}{\left\|x-y\right\|_{2}}\leqslant(1+\epsilon)\sqrt{\frac{M}{N}}$

### v2

@@ C.0.1 Background: standard & modified Meyer’s staircase
    $\begin{split}I\to L^{2}(I)\quad:\quad\mu\mapsto g_{\mu}(t):=\frac{1}{\sqrt[4]{2\pi}\sqrt{\delta}}\exp\left(-\frac{|t-\mu|^{2}}{4\delta^{2}}\right)\end{split}\ \,.$ (10)
@@
    $\begin{split}L^{2}(I)\to\mathbb{R}^{d}\quad:\quad g_{\mu}(t)\mapsto(g_{\mu}(1/d),\dots,g_{\mu}(1))^{\intercal}\end{split}\ \,.$ (11)
@@
    $x_{k}(t)=\frac{1}{\sqrt[4]{2\pi}\sqrt{\delta}}\exp\left(-\frac{|k/d-t|^{2}}{4\delta^{2}}\right)\ \,.$ (12)
@@
    (ii) the local reach of modified Meyer’s staircase is almost the reciprocal of the magnitude of local curvature. The above two aspects indicate that the curve traverses the space with weak self-entanglement. Hence, it suggests the possibility of finding a linear projection $P:\mathbb{R}^{d}\to\mathbb{R}^{d^{\prime}}$ with $d^{\prime}$ much smaller than $d$ such that the projected image $P\gamma$ is a much simpler curve. For example, suppose we perform the linear projection of modified Meyer’s staircase onto its first few principal components: the projected image is a simple curve, and the learning problem can be significantly simplified if we study the regression problem on the projected curve. Hence, to test the performance algorithm for the regression problem, we aim to test curves that are so complex that there is no trivial dimension reduction, e.g., via standard techniques such as Principal Component Analysis.
@@ C.0.2 Meyer helix
    $x_{k}(t)=\frac{1}{\sqrt[4]{2\pi}\sqrt{\delta_{d}}}\cos\left(ak+\frac{t-k/d}{\delta_{d}^{\prime}}\right)G\left(\frac{\left|k/d-t\right|}{\delta_{d}}\right)\ \,,$ (14)
@@
    Moreover, the above properties of complexity are pretty robust: we can also choose Gaussian-type to decay $G(z)=\exp(-z^{2})$ or choose another constant in the variants (e.g., $a,\delta_{d}$, and $\delta_{d}^{\prime}$) and the above properties persist. This indicates that this collection of Meyer helix curves indeed has its complexity growing with ambient dimension $d$. Numerical tests suggest that this collection of curves does not have simple dimension reduction via random projections that preserves geometric properties such that length and reach. Recall the following form of Johnson-Lindenstrauss random projection lemma for manifolds:
@@
    $(1-\epsilon)\sqrt{\frac{M}{N}}\leqslant\frac{\left\|\Phi x-\Phi y\right\|_{2}}{\left\|x-y\right\|_{2}}\leqslant(1+\epsilon)\sqrt{\frac{M}{N}}\ \,.$